### v1

# Can Slow-thinking LLMs Reason Over Time? Empirical Studies in Time Series Forecasting

Abstract
Time series forecasting (TSF) is a fundamental and widely studied task, spanning methods from classical statistical approaches to modern deep learning and multimodal language modeling. Despite their effectiveness, these methods often follow a fast thinking paradigm—emphasizing pattern extraction and direct value mapping, while overlooking explicit reasoning over temporal dynamics and contextual dependencies. Meanwhile, emerging slow-thinking LLMs (e.g., ChatGPT-o1, DeepSeek-R1) have demonstrated impressive multi-step reasoning capabilities across diverse domains, suggesting a new opportunity for reframing TSF as a structured reasoning task. This motivates a key question: Can slow-thinking LLMs effectively reason over temporal patterns to support time series forecasting, even in zero-shot manner? To investigate this, in this paper, we propose TimeReasoner, an extensive empirical study that formulates TSF as a conditional reasoning task. We design a series of prompting strategies to elicit inference-time reasoning from pretrained slow-thinking LLMs and evaluate their performance across diverse TSF benchmarks. Our findings reveal that slow-thinking LLMs exhibit non-trivial zero-shot forecasting capabilities, especially in capturing high-level trends and contextual shifts. While preliminary, our study surfaces important insights into the reasoning behaviors of LLMs in temporal domains—highlighting both their potential and limitations. We hope this work catalyzes further research into reasoning-based forecasting paradigms and paves the way toward more interpretable and generalizable TSF frameworks [^1] https://github.com/realwangjiahao/TimeReasoner.
1 Introduction
Time series forecasting (TSF) is a fundamental machine learning task with wide-ranging applications in finance, energy, healthcare, and transportation box1968some ; deb2017review ; tay2001application . Its core objective is to predict the future trajectory of target variables based on historical observations and auxiliary information—including temporal covariates, static features, and contextual signals hu2016building ; wang2024tabletime . Accurate forecasts are critical for enabling proactive strategies and effective decision-making under uncertainty in real-world scenarios.
Over the past decades, numerous methods have been proposed for TSF lim2021time ; cheng2025comprehensive , which can be broadly categorized into two paradigms: statistical modeling and data-driven learning. Traditional statistical models (e.g., ARIMA zhang2003time , Theta assimakopoulos2000theta , Holt-Winters chatfield1978holt ) assume linearity, stationarity, and seasonality, offering interpretability but limited flexibility in nonstationary settings. In contrast, data-driven approaches, particularly deep learning-based models complex patterns like trends and seasonality via nonlinear architectures, showing strong performance across benchmarks benidis2022deep ; NEURIPS2024_053ee34c ; zhang2024trajectory ; ilbert2024analysing . Recently, with the rapid progress of LLMs, time series forecasting has been explored from a multimodal understanding perspective, leveraging LLMs’ ability to interpret heterogeneous contextual and auxiliary information jiang2024empowering ; zhang2024large . However, most existing attempts adopt LLMs as black-box sequence mappers, without explicitly utilizing their potential for step-wise temporal reasoning, which is crucial for long-horizon and context-aware forecasting tasks tan2024language ; gat2023faithful ; ye2025multisteptimeseriesinference .
Figure 1: Motivation (right): large language models possess both knowledge of time series data (insights into the patterns and forecasting methods) and reasoning capability, which are key factors to build a powerful time series forecaster. Pipeline (left): LLMs use auxiliary information and observed data to take inference-time reasoning, providing interpretable analytical results.
Despite their effectiveness, most TSF methods follow a fast-thinking paradigm jha2016comparative ; lawson2020comparing , mapping historical inputs to future targets via single-step inference with little to no intermediate reasoning chow2024towards . Such models often capture shallow correlations or temporal patterns but struggle with integrating contextual and auxiliary information through deliberate, step-by-step reasoning kauppinen2007modeling . In contrast, slow-thinking LLMs like ChatGPT-o1 and DeepSeek-R1 guo2025deepseek have shown strong multi-step reasoning capabilities in domains such as mathematics, typically by generating intermediate steps before final answers muennighoff2025s1 ; potosnak2025investigating . However, it remains unclear whether such reasoning abilities can be effectively applied to time series forecasting, especially in zero-shot settings without task-specific supervision.
Clearly, these questions are central to rethinking time series forecasting. In this work, we explore whether LLMs—especially those with slow-thinking capabilities—can function as forecasters via reasoning. To this end, we reformulate TSF as a conditional reasoning task, where predictions are guided by historical data, auxiliary variables, and contextual information. Building on this formulation illustrated in Figure 1, we introduce TimeReasoner, a multimodal inference-time reasoning framework that leverages LLMs’ semantic understanding to generate structured forecasts. We further conduct controlled experiments to examine how prompting strategies, temporal setups, and reasoning behavior influence forecasting outcomes.
Q1. Can slow-thinking LLMs perform accurate time series forecasting in zero-shot settings? While recent work has explored LLMs for TSF through prompting, it remains unclear whether slow-thinking LLMs—those capable of multi-step reasoning-can produce reliable forecasts without task-specific training. We evaluate their zero-shot performance against strong supervised baselines.
Q2. How do predict window and lookback window sizes affect LLM-based forecasting? We analyze how varying context lengths influence both forecasting accuracy and the structure of reasoning traces generated by LLMs.
Q3. How robust are LLM-based forecasters to incomplete or noisy input data? We simulate real-world conditions by injecting noise and missing values into the input and examine how LLM-based forecasts degrade under such perturbations.
Q4. Which components of the hybrid prompt are essential for accurate forecasting? We conduct ablation studies on hybrid prompts—including timestamps, context information and raw series-to determine which inputs most affect accuracy and reasoning coherence.
Q5. How do inference-time parameters influence forecasting quality and stability? We evaluate the impact of temperature and model selection on the accuracy of LLM-generated forecasts.
Q6. Do the reasoning traces generated by LLMs offer explainable insights into their predictions? We analyze the intermediate reasoning steps to assess their coherence, alignment with input signals, and potential as a new lens for interpreting model behavior.
2 Preliminaries and Motivation
2.1 Problem Definition
Given a multivariate time series $\mathcal{Y}=\{\mathbf{y}_{t}\}_{t=1}^{T}$, where $\mathbf{y}_{t}\in\mathbb{R}^{d}$ denotes the $d$-dimensional target variables at time step $t$, the goal of multivariate time series forecasting (TSF) is to predict future values $\{\mathbf{y}_{t},\mathbf{y}_{t+1},\ldots,\mathbf{y}_{t+H-1}\}$ over a fixed horizon $H$, based on historical observations and additional context. At each step $t$, the model is provided with a lookback window of target values $\{\mathbf{y}_{t-L},\ldots,\mathbf{y}_{t-1}\}$ of length $L$, and a context vector $\mathcal{Z}\in\mathbb{R}^{k}$ capturing time-invariant metadata, temporal indicators, or semantic descriptors associated with the instance. This formulation highlights the integration of historical dynamics and contextual signals as forecasting conditions.
2.2 Previous Advancements in TSF
TSF has historically evolved along two main modeling paradigms: statistical models cheng2025comprehensive , which rely on handcrafted assumptions like linearity and stationarity, and data-driven models lim2021time ; benidis2022deep , which learn representations from observed data without explicit priors. Classical statistical methods like ARIMA zhang2003time and Holt-Winters chatfield1978holt offer interpretability but struggle with high-dimension or context-rich inputs. In contrast, modern data-driven approaches especially deep learning models like RNNs, TCNs, and Transformers capture complex temporal dependencies and have achieved strong empirical results nie2022time ; liu2023itransformer ; wu2021autoformer ; lin2024cyclenet ; chen2024accelerating . Recently, LLMs have also been explored for TSF, utilizing their capability to process multimodal inputs and understand structured descriptions jin2023time ; gruver2023large .
However, despite architectural differences, most of these methods still follow a fast-thinking paradigm: they generate forecasts in a single forward pass, often without engaging in explicit reasoning over time-evolving context. Yet, many real-world forecasting challenges involve latent dynamics or external conditions like holidays shifting energy usage that require step-wise reasoning over historical trends and auxiliary information. Furthermore, interpretability is essential in high-stakes domains, where models must explain not only what will happen, but why. These challenges motivate a reframing of TSF as a conditional reasoning task, where LLMs are guided via structured prompts to perform multi-step inference. This direction becomes feasible with the emergence of slow-thinking LLMs, which demonstrate strong reasoning capabilities across diverse domains. In this work, we build on this insight and propose TimeReasoner to systematically explore this paradigm.
3 General Framework for Slow-thinking LLMs as TimeReasoner
We use large language models (LLMs) to reformulate the forecasting task as a conditional reasoning problem, where the model generates predictions by interpreting structured natural language prompts. Specifically, at each forecasting step, we construct input patterns that encode the historical target window $\mathcal{Y}$ (conditions) and additional contextual features $\mathcal{Z}$ (descriptors). These components are then filled into a prompt template $T$, which serves as the instruction input to the LLMs.
Under this setting, the LLM is expected to understand the instruction and output the predicted future values through language-based inference. The overall framework of our proposed method TimeReasoner is illustrated in Figure 2, and we describe the detailed prompt construction and reasoning strategies in the subsequent sections.
Hybrid Instructions. To support temporal reasoning, we design multimodal prompts that combine:
  • Raw time series in original space. TimeReasoner takes time series in their original scale as input, without normalization or standardization. Preserving the raw magnitudes allows the model to reason over absolute values and real-world fluctuations that might otherwise be distorted.
  • Timestamp feature. TimeReasoner incorporates original timestamps into the prompt, enabling it to recognize both absolute temporal positions (e.g., specific hours or dates) and relative patterns (e.g., periodicity or time intervals), which are crucial for temporal reasoning.
  • Contextual feature descriptions. TimeReasoner is provided with natural language descriptions that offer contextual information like domain-level knowledge and channel-specific semantics. These help align numerical data with real-world meaning, enhancing interpretability and task relevance.
Figure 2: An overview of our proposed TimeReasoner framework. Left: hybrid instructions, some necessary auxiliary information to assist LLMs in reasoning. Right: reasoning and forecasting, TimeReasoner’s slow thinking reasoning steps.
Reasoning Strategies. We explore three distinct reasoning paradigms to evaluate the capabilities of slow-thinking LLMs (see Appendix C for more details and the prompt template):
  • One-Shot Reasoning. In one-shot multi-step reasoning, TimeReasoner performs a single, comprehensive reasoning pass on the given input. During one attempt, it processes all necessary information and internally executes the multi-step logic required. The model then immediately outputs the complete set of predicted results , without any subsequent iterative.
  • Decoupled Reasoning. For decoupled multi-step reasoning, TimeReasoner aims to generate all results as a single, comprehensive output. Unlike one-shot reasoning, its process is decoupled: the model generates a partial thought, reflects on it, and then continues, forming a generate–reflect–generate loop within a single inference, enabling refined multi-step reasoning.
  • RollOut Reasoning. In rollout multi-step reasoning, TimeReasoner generates predictions incrementally, producing one part of the result in each step. This process involves several iterative rounds, where each output serves as context for the next, enabling TimeReasoner to gradually build toward a comprehensive final result.
4 Experimental Setup
Datasets. To cover a broader range of practical scenarios, we utilize datasets from various domains. These datasets include: the ETT dataset zhou2021informer which records electricity load from 2016 to 2018, Exchange lai2018modeling which tracks foreign exchange rates from 1990 to 2016, Wind lai2018modeling containing wind measurements from 2020 to 2021, AQ zhang2017cautionary providing four years of air quality data, NASDAQ feng2019temporal which offers comprehensive stock market time series including opening and closing prices, trading volumes, and daily high-low values and VitalDB from which we select a single series originally sampled at 3-second intervals cheng2024hmf ; lee2022vitaldb . Following standard protocol, each dataset is chronologically split into training, validation, and testing sets. Further details are provided in Appendix B.1.
Baselines. Our baseline methods include the following competitive methods: deep learning-based methods: Autoformer wu2021autoformer , FEDformer zhou2022fedformer , PatchTST nie2022time , iTransformer liu2023itransformer , TimeXer wang2024timexer and DLinear zeng2023transformers . For LLM-based methods, we select GPT4TS zhou2023one , Time-LLM jin2023time and LLMTime gruver2023large . Detail information ca be found in Appendix B.2.
Implement Details. We choose DeepSeek-R1 as the base model for TimeReasoner due to its open-source nature and transparent chain of thought. Inference is performed via the API  [^2] https://www.deepseek.com/, with temperature, top-p, and max-tokens set to 0.6, 0.7, and 4,096. For each setting, three independent experiments are conducted, and we report the average. For baseline methods, we use ADAM kingma2014adam as the optimizer and all experiments are implemented in PyTorch paszke2019pytorch and run on a single NVIDIA RTX 4090D 24GB GPU with a fixed random seed. Baseline code is available at  [^3] https://github.com/thuml/Time-Series-Library. Lookback windows are $T$ = 36 for NASDAQ, 300 for VitalDB, and 96 for other datasets; predict windows are $H$ = 36 for NASDAQ, 200 for VitalDB, and 96 for others. For each dataset, we use a channel-independent method by averaging results from random test samples.
Evaluation. We evaluate forecasting performance using two standard regression metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). MSE captures the average squared difference between predictions and ground truth, while MAE reflects the average prediction error. For each dataset, the last segment is used for testing, with earlier parts for training and validation. Results are reported across multiple forecast horizons to assess both short- and long-term accuracy.
5 Benchmarking Time Series Reasoning Methods (Q1)
Performance Comparison with Baseline Methods. Our multivariate forecasting results, detailed in Table 1, demonstrate TimeReasoner’s outstanding performance over other baseline models. Results on ten datasets show that TimeReasoner always achieves significant forecasting results. Especially on the ETTh1 and AQShunyi, the forecasting values are advantageous compared to the baseline method.
Forecasting Performance on Complex Datasets. The surprising finding is that TimeReasoner particularly excels on datasets characterized by complex temporal dynamics. On datasets like AQWan, AQShunyi and Wind, TimeReasoner consistently surpasses other methods by effectively capturing underlying dependencies. These findings underscore TimeReasoner’s robust capability in tackling difficult forecasting tasks and its generalization to complex time series. It cannot be ignored that performance remains competitive with baselines on simpler datasets, showing no substantial advantage. This suggests that while the model excels in modeling complex temporal dynamics, its performance on less intricate time series remains robust yet aligns with established benchmarks.
Table 1: Performance comparison of TimeReasoner and baseline models with best values in bold and second-best underlined. MSE $\downarrow$ and MAE $\downarrow$ are used as the evaluation metrics. Overall, TimeReasoner achieves superior performance across datasets.
Methods Metric ETTh1 ETTh2 ETTm1 ETTm2 Exchange AQWan AQShunyi Wind NASDAQ VitalDB
AutoFormer MSE 7.4310 13.4821 15.3155 7.0926 0.0012 19662.2860 23335.7346 1659.3374 0.0008 76.7812
MSE 1.5719 2.2928 2.0689 1.7454 0.0227 47.7715 50.7481 19.3377 0.0226 7.2136
FEDFormer MSE 16.1369 13.0165 22.1313 19.1682 0.0013 29623.8156 19774.4204 2168.1601 0.0009 82.3565
MSE 2.2300 2.3180 2.4822 2.7404 0.0244 55.5783 51.1552 23.9338 0.0215 7.6731
PatchTST MSE 9.2430 10.9489 16.3988 5.6375 0.0009 12528.5571 16747.3794 2013.9736 0.0007 50.7402
MSE 1.6500 1.9930 1.9294 1.3834 0.0198 39.8232 42.0165 20.1837 0.0211 5.6983
iTransformer MSE 7.5195 9.9716 12.7432 5.7392 0.0010 13577.9277 17654.6335 1600.0376 0.0008 79.2864
MSE 1.4983 1.9010 1.6374 1.4422 0.0202 39.8069 41.8849 17.9244 0.0235 7.3322
TimeXer MSE 8.4878 11.4123 14.0473 5.5801 0.0009 14509.0913 16505.3908 1668.6824 0.0007 65.8567
MSE 1.5430 2.0606 1.7534 1.4238 0.0193 41.2693 40.9784 18.2386 0.0230 6.1253
DLinear MSE 7.7656 10.3584 14.0352 7.8604 0.0014 21090.0523 21077.8247 1617.3343 0.0007 68.2254
MSE 1.4981 2.0567 1.7591 1.8981 0.0253 51.1701 50.5644 18.3691 0.0215 6.7148
GPT4TS MSE 6.9454 9.7156 15.9028 5.6327 0.0009 13543.6153 16821.9321 1786.3085 0.0010 65.3715
MSE 1.4127 1.8891 1.9364 1.4743 0.0200 39.7527 41.8708 18.2427 0.0253 6.7207
Time-LLM MSE 6.8152 9.9876 15.8078 5.6787 0.0010 13405.9845 16678.0982 1773.9651 0.0011 65.9761
MSE 1.4115 1.8880 1.9350 1.4733 0.0199 39.7481 41.8598 18.2365 0.0252 6.7151
LLMTime MSE 10.6261 11.2011 14.6394 6.9001 0.0026 29141.5143 30266.3043 3979.7996 0.0021 99.5761
MSE 1.6987 1.5876 1.9543 1.5109 0.0345 65.6789 60.3456 30.1234 0.0234 8.9054
TimeReasoner MSE 5.4469 8.6020 14.4327 6.4384 0.0009 11305.6345 12874.9412 1556.5316 0.0008 79.4890
MSE 1.1984 1.6811 1.7179 1.4209 0.0168 36.1614 37.0193 17.5034 0.0215 6.9735
6 Evaluation of TimeReasoner in Different TSF Settings (Q2)
6.1 Performance Comparison w.r.t Varying Predict Window
In previous approaches, forecasting errors generally exhibit a positive correlation with the predict window length $H$ since longer horizons introduce greater error propagation. Whether this principle still holds for TimeReasoner, which rely on step-by-step slow reasoning rather than learned representations, remains unclear. To investigate this, we fix the lookback window $L=96$ and gradually increase the predict window $H=\left\{48,96,144,192\right\}$. As shown in Figure 3, MSE increases with a longer predict window. This also proves that the previous principle is applicable to TimeReasoner. This performance degradation suggests that TimeReasoner’s reasoning becomes more difficult since it extrapolates further from observed data, limiting its effectiveness for long-term forecasting.
6.2 Performance Comparison w.r.t Varying Lookback Window
Like the principle elaborated in 6.1, the length of the lookback window $L$ determines how much historical context one model receives, and previous methods benefit monotonically from longer histories. For slow thinking-based TimeReasoner, we may wonder whether the longer the lookback window, the more beneficial it is for LLMs. To discovery this unknown, we fix the predict window $H=96$ and vary the lookback window $L=\left\{48,96,144,192\right\}$. As shown in Figure 3, we observe a nonmonotonic relationship between MSE and $L$: initially, increasing $L$ improves performance as the model gains access to richer patterns and contextual cues. However, beyond a certain point, more history begins to degrade performance, likely due to the irrelevant or noisy data. This suggests that while TimeReasoner can leverage historical context to enhance reasoning, overly long inputs may dilute relevant signals and hinder generalization. Striking a balance between context richness and conciseness is thus crucial for optimal forecasting.
Figure 3: Left: The positive correlation between MSE and the length of predict window with experiments on ETTh2, Exchange datasets. Right: The relationship between MSE and the lookback window is not proportional with experiments on AQWan, Wind datasets.
6.3 Robustness Evaluation w.r.t. Missing Input (Q3)
Missing values frequently occur in real-world time series due to issues like sensor failures che2018recurrent . While traditional models depend on imputation or data cleaning, reasoning-based models such as TimeReasoner can process raw, incomplete inputs. We evaluate the model’s robustness to missing inputs by comparing its performance across four input scenarios: (1) complete data with no missing values, (2) missing time series values without imputation (WI), (3) time series with missing values imputed using None placeholder (NPI), and (4) missing values using linear interpolation for imputation (LII) chow1971best . For a detailed description, please refer to Appdedix B.3.
Table 2: Comparison with different missing data handling strategies. WI refers to no imputation, NPI means that missing values are marked as None placehold, and LII refers to linear interpolation. This demonstrates the robustness of TimeReasoner to missing values and the importance of preserving the full temporal structure.
Models ETTh1 ETTm1
MSE MAE MSE MAE
TimeReasoner 5.4469 1.1984 14.2055 1.6984
w/ WI 9.4523 1.4781 24.4894 2.2308
w/ NPI 6.6059 1.2498 14.5843 1.7418
w/ LII 5.7765 1.2218 14.4251 1.7154
As shown in Table 2, TimeReasoner achieves the best performance with fully observed data, highlighting its capacity to capture comprehensive temporal dependencies. When linear interpolation is used to fill missing values, the model still achieves outstanding prediction results, highlighting the robustness of the TimeReasoner model. Imputing missing values with None results in moderate performance degradation, as the model must contend with increased uncertainty during inference. In contrast, removing missing entries entirely leads to a more substantial decline with decrement of 72.39% on ETTm1 dataset, indicating that the disruption of temporal structure significantly impairs the model’s reasoning. These findings suggest that, while TimeReasoner exhibits a degree of robustness to missing data, preserving the temporal structure, even with placeholders is preferable to deletion.
7 Effects of Key Settings in Hybrid Instruction (Q4)
Table 3: Ablation study on timestamps, context information and normalization strategies.
Key Settings Models ETTm1 Wind NASDAQ
MSE MAE MSE MAE MSE MAE
Components TimeReasoner 5.4469 1.1984 1556.5316 17.5034 0.0008 0.0215
Timestamps w/ removing timestamps 25.3058 2.4705 3278.3356 29.7851 0.0021 0.0476
w/ forward shifting 7.9222 1.3044 1570.5614 17.8587 0.0009 0.0218
w/ backward shifting 6.4596 1.2640 2195.4411 19.3174 0.0010 0.0218
Context Information w/o context information 7.8781 1.4325 2239.5041 19.4609 0.0009 0.0238
Normalization w Z-score Normalization 8.7691 1.4601 2076.8085 18.8674 0.0011 0.0236
w RevIN 122.0912 11.7891 6641.7854 67.1268 0.0025 0.0671
7.1 The Effect of Timestamps on TimeReasoner
Timestamps play a vital role in time series forecasting by serving as positional encodings that capture absolute and relative temporal information wang2024rethinking ; zeng2024much . For TimeReasoner, it remains unclear whether its performance relies on timestamps, and if so, whether it exploits absolute clock time or relative intervals. To investigate this, we design two perturbation strategies: removing, which discards all timestamps and treats the series as unordered; and shifting, where a time-shift operator $\mathcal{T}_{k}$ uniformly offsets all timestamps by a fixed step $k$. Full details are provided in Appendix B.4.
As shown in Table 3, applying a forward or backward timestamp shift leads to mild performance degradation on all three datasets, suggesting that even slight temporal distortions can affect zero-shot forecasting. On Wind dataset, the backward shift causes a more noticeable increase in MSE, indicating dataset-specific sensitivity. Most notably, removing timestamps results in substantial drops in both MSE and MAE, highlighting the critical role of temporal information. These results show that while TimeReasoner is somewhat robust to minor perturbations, it fundamentally relies on both absolute and relative timestamps for accurate forecasting.
7.2 The Analysis of Contextual Information in Prompt of TimeReasoner
Incorporating contextual information—auxiliary metadata tied to the dataset—has gained attention in time series modeling yamaguchi2025citras ; lim2021temporal . TimeReasoner integrates such context into prompts, and an ablation study is conducted to assess its impact on forecasting accuracy. As shown in Table 3, contextual input significantly improves performance on the Wind dataset by reducing both MSE and MAE, suggesting that semantic cues like location or seasonality enhance predictions. However, on the NASDAQ dataset, its effect is minimal, likely due to high volatility and latent, unobserved factors. These results highlight the domain-dependent utility of contextual information in time series forecasting.
7.3 The Impact of Normalization for TimeReasoner
Normalization is widely used in time series forecasting to accelerate model convergence and reduce distribution drift liu2023adaptive . However, its effectiveness for TimeReasoner remains unclear. To investigate this, we compare two popular normalization strategies: Z-score normalization patro2015normalization and RevIN kim2021reversible . Detailed descriptions of both methods are provided in Appendix B.5. As shown in Table 3, TimeReasoner achieves better performance on the unnormalized time series data compared to when either Z-score or RevIN normalization is applied. This result suggests that the model benefits from directly leveraging raw temporal patterns and intrinsic data characteristics, which may be distorted by statistical normalization. While normalization typically stabilizes training and reduces scale sensitivity, it can obscure critical features such as long-term trends, periodicity, and abrupt transitions. These findings highlight that, in some forecasting scenarios, preserving the original data distribution may enhance model reasoning by retaining richer temporal semantics.
8 Evaluation of Various Reasoning Strategies
One hypothesis is that RollOut Reasoning should be more powerful in short-term forecasting because its multi-round reasoning mechanism will have lower cumulative errors than long-term forecasting. In terms of long-term forecasting, Decoupled Reasoning should have more advantages because it emphasizes the reasoning optimization of LLMs in one reasoning. We fix the lookback window $L=96$ and compare the of performance under four predict window lengths $H=\left\{48,96,144,192\right\}$. As shown in the Figure 4, One-Shot Reasoning performs well across all predict windows. RollOut Reasoning achieves the best performance at shorter windows (48 and 96). However, as the prediction window extends to 144 and 192, its error increases more rapidly, and Decoupled Reasoning surpasses it to become the best-performing method. This suggests that while One-Shot may struggle with long-range dependencies and RollOut suffers from error accumulation, Decoupled Reasoning benefits from internal refinement, maintaining stability and higher accuracy over longer sequences.
Figure 4: Comparison of TimeReasoner’s forecasting capabilities under the three reasoning strategies with experiments conducted on ETTh2 and Exchange datasets.
9 Hyper-paramter Sensitivity Analysis (Q5)
9.1 Performance Comparison under Varying Temperatures
Temperature controls the level of uncertainty in LLM outputs during generation. To explore how temperature affects TimeReasoner’s reasoning behavior, we conduct a series of controlled experiments. As show in Figure 5, performance exhibits a non-linear relationship with temperature: as temperature increases from 0, the MSE initially decreases, reaching its minimum at 0.6, and then rises again. Possible reasons are that low temperatures yield stable but rigid forecasts, leading missing nuanced temporal variations. High temperatures result in high-variance predictions that degrade reliability. The optimal performance at a moderate temperature suggests that a balanced degree of uncertainty enhances the model’s reasoning ability.
Figure 5: Left: As the temperature increases, MSE and MAE first decrease and then increase. 0.6 is a more suitable parameter. Right: Comparison of TimeReasoner’s forecasting capabilities when utilizing different base LLMs. DeepSeek-R1 demonstrates the best results.
9.2 Performance Comparison across Base LLMs
In this section, we compare the performance of TimeReasoner when utilizing various base reasoning LLMs. Specifically, we select five representatives, namely DeepSeek-R1, GPT-o1 mini, Gemini-2.5 Pro, QwQ-plus, and DeepSeek-R1-Distill-Llama-70B. As shown in Figure 5, DeepSeek-R1 always demonstrates the best performance while DeepSeek-R1-Distill-Llama-70B showed the weakest performance across both evaluations.
10 Analysis of COT in TimeReasoner & Case Study Analysis (Q6)
10.1 Empirical Analysis of COT
Reasoning Patterns in TimeReasoner. Building on TimeReasoner’s strong zero-shot forecasting capabilities, we analyze its internal reasoning process, which qualitatively consists of three stages: (1) Data analysis and pattern recognition, where the model identifies structural features such as intervals, frequency, trends, and seasonality; (2) Forecasting strategy evaluation, where it weighs methods ranging from seasonal baselines to trend extrapolation, reflecting on their trade-offs; and (3) Reflection and contextual assessment, where the model critiques its own reasoning, noting strengths like capturing seasonality and limitations such as lacking complex statistical mechanisms. This process demonstrates adaptive and context-aware thinking.
In-depth Analysis of Chain-of-Thought. An intuitive guess is that the lower the ADF value of the input time series sample, the lower its reasoning cost, the lower of token number of chain of thought. As shown in Figure 6, the relationship between the token and the ADF of the input sample. From the scatter, we can see that there is no direct relationship between the two.
We interpret the number of tokens used during inference as an indicator of the model’s reasoning depth. To explore its impact, we categorized the results into low, medium, and high MSE groups under three different token length settings: small, medium, and large since we have conducted three runs on ome sample. As shown in Figure 6, our analysis shows that higher token counts consistently correspond to a greater proportion of samples achieving the lowest MSE. Conversely, lower token counts are associated with poorer performance. This pattern aligns with our intuition: increased token usage allows the model to perform deeper reasoning, thereby enhancing its ability to solve forecasting tasks effectively.
Figure 6: Left: The relationship between the token number of chain of thought and ADF test value. Right: The proportion of high, medium, and low MSEs caused by changes in token length in three independent experiments on a single sample.
Figure 7: A case study of how TimeReasoner reason 96-96 time series forecasting. This example shows how TimeReasoner apply slow thinking to this task.
10.2 Case Study Analysis
Figure 7 shows one case of inference time reasoning of TimeReasoner. We can clearly see that LLM imitates human experts to take slow-thinking reasoning. It first analyzes the pattern of the data, then conducts a detailed analysis, and then determines the model to be used, and finally obtains the forecasting result. This reflects the careful consideration of the model when making inferences.
11 Conclusion and Future Work
In this work, we investigated whether large language models (LLMs)—particularly those with slow-thinking capabilities—can reason over temporal patterns to support time series forecasting. We reformulated the forecasting task as a conditional reasoning problem, where structured prompts encode historical observations and contextual features to guide LLMs in generating future trajectories. To support this paradigm, we introduced TimeReasoner, a prompting-based framework that incorporates hybrid inputs—including raw time series, timestamps, and semantic descriptors—and supports multiple inference-time reasoning strategies such as one-shot, decoupled, and roll-out reasoning. Experiments across diverse TSF benchmarks show that slow-thinking LLMs exhibit strong zero-shot forecasting performance. More importantly, their intermediate reasoning traces offer interpretable insights into the temporal dynamics underlying the data.
We acknowledge that this is an early step toward building reasoning-centric, multimodal forecasting systems. In future work, we aim to expand TimeReasoner to more complex real-world settings and investigate more principled reasoning mechanisms to further enhance the reliability, interpretability, and generalization of LLM-based time series forecasters.
Appendix
Appendix A Related Work
A.1 Time Series Forecasting
Time series forecasting has been extensively studied over the past few decades cheng2025comprehensive . Early methods, such as ARIMA box1968some ; zhang2003time , employed an auto-regressive approach and forecasted using a moving average framework. While theoretical guarantees were established, these traditional methods often require data with ideal properties, which do not always align with real-world scenarios hewamalage2021recurrent . With the growing availability of data and advances in computing power, numerous deep learning-based models have emerged, typically following a sequence-to-sequence paradigm salinas2020deepar . Recurrent neural networks were among the first to capture temporal dependencies by summarizing past information in time series data wen2017multi . However, these architectures suffer from limitations such as a restricted receptive field and error accumulation due to recursive inference lai2018modeling , which undermine forecasting accuracy. To improve performance, advanced architectures have been developed to capture long-range dependencies zhou2021informer , including those with self-attention mechanisms and convolutional networks li2019enhancing ; liu2022scinet . Recent work has also integrated traditional techniques, like trend-seasonal decomposition and time-frequency transformations zhou2022fedformer ; liu2024generative , into neural networks to better capture time series data characteristics wen2020fast ; zhou2022film ; cheng2025convtimenet . Additionally, a recent study showed that even a simple linear network enhanced with decomposition can achieve competitive performance zeng2023transformers . Furthermore, slice-based methods have demonstrated superior accuracy in long-term forecasting tasks nie2022time ; zhang2023crossformer ; cheng2025instructime .
A.2 Large Language Models for Time Series
In recent years, large language models (LLMs) have garnered significant attention for their remarkable ability to understand and generate human-like text kumar2024large ; yao2024survey . While traditionally applied to natural language processing tasks zhang2024large , these models have also shown great potential in time series analysis  jin2024position . The application of LLMs to time series can be broadly categorized into two primary paradigms: fine-tuning and prompt-based zero-shot learning jiang2024empowering . In fine-tuning, a pre-trained LLM is further trained on time series data specific to a given task chang2023llm4ts ; chang2024align , enabling it to learn task-specific patterns and improve performance  liu2025calf ; jin2023time . However, this approach often requires substantial labeled data and computational resources cheng2025comprehensive . On the other hand, prompt-based zero-shot methods leverage the model’s pre-existing knowledge gruver2023large , guiding its predictions through well-crafted prompts without the need for task-specific training liu2024lstprompt . Although zero-shot methods are more flexible and resource-efficient tang2025time , they may not always achieve the same level of performance as fine-tuned models, especially on specialized time series tasks wang2024tabletime . Both paradigms highlight the growing interest in applying LLMs to time series, although challenges remain in optimizing effectiveness for these tasks.
A.3 Inference-Time Reasoning
Inference-Time Reasoning zeng2025revisiting refers to the strategy of enhancing a pre-trained model’s performance on complex reasoning tasks by allocating additional computational resources during the inference phase singhi2025solve ; yang2025towards . These resources may include extended inference time, more computational steps, or advanced reasoning strategies such as multi-path sampling and iterative refinement leehanchung . Unlike traditional training-time scaling, which focuses on improving a model’s capabilities during training through increased parameters, data, or computation parashar2025inference , inference-time reasoning utilizes a larger computational budget at inference to unlock deeper reasoning potential within the model—potential zhang2025and that might otherwise remain dormant during rapid inference. This approach improves performance without altering the model’s weights chen2024simple .
Several cutting-edge large language models demonstrate the effectiveness of inference time reasoning sui2025meta . For instance, DeepSeek-R1 guo2025deepseek can benefit from techniques such as iterative refinement or self-correction, utilizing additional computational resources during inference to enhance performance on complex tasks, leveraging its strengths in coding and reasoning mercer2025brief . Similarly, models such as Gemini-2.5-pro team2023gemini are expected to excel in sophisticated reasoning tasks when granted extended processing time or multi-step inference protocols enabled by inference-time reasoning. OpenAI’s models, including versions such as GPT-o1 hasei2025comparative , show significant improvements on challenging reasoning problems when employing techniques such as chain-of-thought prompting or step-by-step decomposition, both of which increase the computational budget during inference wei2022chain . These examples highlight the growing importance of inference-time reasoning in pushing the boundaries of AI reasoning by trading inference-time computation for higher accuracy muennighoff2025s1 .
Appendix B Experimental Details
B.1 Datasets
Here is a detailed introduction to the datasets: (1) ETT records the oil temperature and load features of the 139 electricity transformers from July 2016 to July 2018. It is made up of 4 sub datasets where ETThs are 140 sampled per hour and ETTms are sampled every 15 minutes. (2) Exchange collects the daily 141 exchange rates of 8 countries from 1990 to 2016. (3) Wind collects records from 2020-2021 at 142 15-minute intervals. (4) AQ records the air quality from two measurement stations over a period of 4 years. It consists of two sub datasets and two datasets are sampled per hour. (5) NASDAQ records opening price, closing price, trading volume, lowest price, and highest price of nasdaq stock market. (6) VitalDB initially includes 6,388 records, resampled at 3-second intervals. We choose one record to conduct our experiment. Table 4 provides more information, including timesteps, sampling frequency, etc.
Table 4: Statistics information of experimental datasets.
Dataset Domain Timestamps Features Frequency
ETTh1 & ETTh2 Electricity 17,420 7 1 hour
ETTm1 & ETTm2 Electricity 69,680 7 15 mins
AQWan & AQShunyi Environment 35,064 11 1 hour
Exchange Economy 7,588 8 1 day
NASDAQ Economy 1,244 5 1 day
Wind Energy 48,673 7 15 mins
VitalDB Healthcare 2,855 2 3s
B.2 Baselines
  • Autoformer wu2021autoformer : Introduces a decomposition-based Transformer architecture with an auto-correlation mechanism to capture long-term dependencies in time series forecasting.
  • FEDformer zhou2022fedformer : Combines seasonal-trend decomposition with frequency domain analysis using Fourier transforms to enhance long-term time series forecasting.
  • PatchTST nie2022time : Segments time series into patches and applies a channel-independent Transformer to improve long-term forecasting accuracy.
  • iTransformer liu2023itransformer : Applies attention mechanisms on inverted dimensions, embedding time points as variate tokens to capture multivariate correlations in time series data.
  • TimeXer wang2024timexer : Integrates exogenous variables into Transformer-based time series forecasting, enhancing predictions by incorporating external information.
  • DLinear zeng2023transformers : A simple and efficient linear model that decomposes time series into trend and seasonal components, achieving competitive performance with reduced complexity.
  • GPT4TS zhou2023one : Utilizes pre-trained GPT-2 models for time series forecasting by feeding time series patches in a channel-independent manner.
  • Time-LLM jin2023time : Reprograms large language models for time series forecasting by converting time series data into textual prompts, enabling few-shot and zero-shot learning capabilities.
  • LLMTime gruver2023large : Employs large language models for zero-shot time series forecasting by encoding numerical data as text and generating predictions through text completions.
B.3 Robustness Evaluation
Missing Value Construction. Given an input time series $\mathcal{Y}_{1:T}\in\mathbb{R}^{T\times d}$, we randomly drop values to simulate missing data. Let $\mathbf{M}\in\{0,1\}^{T\times d}$ be a binary mask where each entry is sampled independently from a Bernoulli distribution:
$\mathbf{M}_{t,i}\sim\mathrm{Bernoulli}(1-p),$ (1)
where $p$ is the missing rate. The corrupted input with missing values is then:
$\tilde{\mathcal{Y}}_{t,i}=\begin{cases}\mathcal{Y}_{t,i},&\text{if }\mathbf{M}_{t,i}=1\\ \texttt{NaN},&\text{if }\mathbf{M}_{t,i}=0\end{cases}$ (2)
Missing Value Handling Methods. We compare three methods for handling missing values in $\tilde{\mathcal{Y}}_{1:T}$:
  • WI: Without Imputation. No modification is applied. The input remains: $\tilde{\mathcal{Y}}_{1:T}^{\text{WI}}=\tilde{\mathcal{Y}}_{1:T}$.
  • NPI: None Placeholder Imputation. Missing entries are replaced with a placeholder None:
    $\tilde{\mathcal{Y}}_{t,i}^{\text{NPI}}=\begin{cases}\mathcal{Y}_{t,i},&\text{if }\mathbf{M}_{t,i}=1\\ \texttt{None},&\text{if }\mathbf{M}_{t,i}=0\end{cases}$ (3)
  • LII: Linear Interpolation Imputation. Missing values are filled by linear interpolation along the temporal dimension. For each missing $\mathcal{Y}_{t,i}$, we estimate:
    $\tilde{\mathcal{Y}}_{t,i}^{\text{LII}}=\mathcal{Y}_{t_{1},i}+\frac{t-t_{1}}{t_{2}-t_{1}}\left(\mathcal{Y}_{t_{2},i}-\mathcal{Y}_{t_{1},i}\right),\quad\text{for }t_{1}<t<t_{2},$ (4)
    where $t_{1}$ and $t_{2}$ are the nearest previous and next time points where the value is observed.
B.4 Timestamp Perturbation Methods
To investigate reliance on temporal information of TimeReasoner model, we apply two perturbation strategies to the timestamp component of the input sequence: removing and shifting timestamps. These modifications allow us to assess whether the model is performing genuine temporal reasoning or merely pattern matching based on implicit position encoding.
Removing. In the removing setting, we discard all timestamp information and present the model with only the sequence of feature vectors: $\{y_{1},y_{2},\dots,y_{T}\}$ effectively treating the input as an unordered set of observations. This tests the model’s ability to forecast without access to any explicit temporal cues or ordering. The assumption here is that if the model still performs well, it might be relying on content-based pattern recognition rather than learning from temporal progression.
Shifting. In the shifting setting, we perturb the original timestamps by applying temporal shift operators. Given a sequence $\mathcal{Y}=\{(t_{1},y_{1}),(t_{2},y_{2}),\dots,(t_{T},y_{T})\}$, we define two shift operations:
$\mathcal{T}_{k,\text{forward}}(\mathcal{Y})=\{(t_{1}+k,y_{1}),(t_{2}+k,y_{2}),\dots,(t_{T}+k,y_{T})\},$ (5)
$\mathcal{T}_{k,\text{backward}}(\mathcal{Y})=\{(t_{1}-k,y_{1}),(t_{2}-k,y_{2}),\dots,(t_{T}-k,y_{T})\},$ (6)
where $k\in\mathbb{Z}_{>0}$ denotes the magnitude of the temporal shift.
B.5 Normalization Strategies
Normalization is a key preprocessing step in time series forecasting. We compare two strategies: reversible instance normalization (RevIN) kim2021reversible and Z-score normalization, which differ in how they compute and apply normalization statistics.
Reversible Instance Normalization. Given an input time series $\mathcal{Y}_{1:T}\in\mathbb{R}^{T\times d}$, RevIN computes the mean $\mu\in\mathbb{R}^{d}$ and standard deviation $\sigma\in\mathbb{R}^{d}$ along the temporal dimension for each feature. Then the input time series is normalized as:
$\mathcal{Y}_{1:T}=\frac{\mathcal{Y}_{1:T}-\mu}{\sigma},$ (7)
where $\mu_{d}=\frac{1}{T}\sum_{t=1}^{T}y_{t,d}$ and $\sigma_{d}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}(y_{t,d}-\mu_{c})^{2}+\epsilon}$. This per-instance normalization adapts to non-stationary distributions and allows test-time statistics to differ from training.
Z-score Normalization. Z-score normalization, by contrast, computes global statistics $\mu_{\text{train}},\sigma_{\text{train}}\in\mathbb{R}^{d}$ from the training dataset. These are calculated over all training sequences and time steps, and then applied uniformly:
$\mathcal{Y}_{1:T}=\frac{\mathcal{Y}_{1:T}-\mu_{\text{train}}}{\sigma_{\text{train}}},$ (8)
Appendix C Reasoning Strategies and Prompt Details
Following the notations given in 2.1, we will introduce the formal definitions of three reasoning strategies and the corresponding template of prompt.
C.1 One-Shot Reasoning
One-Shot Reasoning refers to getting the final answer through a one-time iteration with LLMs. The formula can be described as follows:
$\{{\mathbf{\hat{y}}_{t+i}\}^{H}_{i=1}}=f(\{\mathbf{y}_{t+i}\}^{0}_{i=-L+1},\mathcal{Z}),$ (9)
where $f(·)$ denotes a reasoning LLM, $\mathcal{Z}\in\mathbb{R}^{k}$ denotes a context vector capturing time-invariant metadata, temporal indicators, or semantic descriptors associated with the instance. The following LABEL:reasoning1 is the template of this reasoning strategy.
Table 5: Template of One-Shot Reasoning strategy. Contents Enclosed in {} Will Be Replaced with Specific Information, Including Dataset Description, Channel Information, and Historical Data.
Here is the {Channel Information} data of the {Dataset Name} dataset. Here is the dataset description: {Dataset Description}. I will now give you data for the past 96 recorded dates, and please help me forecast the data for the next 96 recorded dates. The data is as follows: {Historical Time Series Data}. Please give me the complete data for the next 96 recorded dates. You can output the answer inside <answer></answer>.
C.2 Decoupled Reasoning
To improve forecasting accuracy over long horizons, we adopt a segment-wise generation strategy that decomposes the overall forecast window into shorter sub-windows. Specifically, given a total forecast horizon of length $H$, we divide it into smaller predict windows: $\{H_{1},H_{2},...,H_{n}\}$ Instead of requiring LLMs to continuously give all predictions in one generation, we encourage LLMs to think multiple times in one generation, and give partial predictions after each thought until the full prediction is given. Note that the entire sequence is still generated in a single forward pass.
$\{{\mathbf{\hat{y}}_{t+i}\}^{H}_{i=1}}=[\{{\mathbf{\hat{y}}_{t+i}\}^{H_{1}}_{i=1}},\{{\mathbf{\hat{y}}_{t+i}\}^{H_{2}}_{i=H_{1}+1}},...,\{{\mathbf{\hat{y}}_{t+i}\}^{H_{n}}_{i=H_{n-1}+1}}],$ (10)
Table 6: Template of Decoupled Reasoning. Contents Enclosed in {} Will Be Replaced with Specific Information, including Dataset Description, Channel Information, and Historical Data.
Here is the {Channel Information} data of the {Dataset Name} dataset. The {Dataset Name} dataset is {Dataset Description}. I will now give you data for the past 96 recorded dates, and please help me forecast the data for the next 96 recorded dates. The data is as follows: {Historical Time Series Data}. Please give me the complete data for the next 96 recorded dates. You should break down the task into 4 steps. For each short-term prediction, please focus on trends of the input sequence. After one step generation, please take a reflection on your generation. You can output the answer inside <answer></answer>.
C.3 RollOut Reasoning
Unlike Decoupled Reasoning, RollOut Reasoning uses a multi-round dialogue mechanism to make predictions. Specifically, supposing that we break down the predict window $H$ into smaller predict windows: $\{H_{1},H_{2},...,H_{n}\}$. We start with forecasting $H_{1}$:
$\{{\mathbf{\hat{y}}_{t+i}\}^{H_{1}}_{i=1}}=f(\{\mathbf{y}_{t+i}\}^{0}_{i=-L+1},\mathcal{Z}),$ (11)
and then we loop the following task based the forecasting result, note that the {Historical Time Series Data} in the i-th iteration includes the previously forecasting values::
$\{{\mathbf{\hat{y}}_{t+i}\}^{H_{2}}_{i=H_{1}+1}}=f(\{\mathbf{y}_{t+i}\}^{-1}_{i=-L},\{{\mathbf{\hat{y}}_{t+i}\}^{H_{1}}_{i=1}},\mathcal{Z}),$ (12)
We iterate this process until we get all the prediction results:
$\{{\mathbf{\hat{y}}_{t+i}\}^{H}_{i=1}}=[\{{\mathbf{\hat{y}}_{t+i}\}^{H_{1}}_{i=1}},\{{\mathbf{\hat{y}}_{t+i}\}^{H_{2}}_{i=H_{1}+1}},...,\{{\mathbf{\hat{y}}_{t+i}\}^{H_{n}}_{i=H_{n-1}+1}}],$ (13)
Table 7: Template of RollOut Reasoning (i-th iteration). Contents Enclosed in {} Will Be Replaced with Specific Information, including Dataset Description, Channel Information, and Historical Data.
Here is the {Channel Information} data of the {Dataset Name} dataset. The {Dataset Name} dataset is {Dataset Description}. I will now give you data for the past {Lookback Window+$\sum_{k=1}^{i-1}H_{k}$} recorded dates, and please help me forecast the data for the next {Predict Window ($H_{i}$)} recorded dates. The data is as follows: {Historical Time Series Data}.
Please give me the complete data for the next {Predict Window ($H_{i}$)} recorded dates. You can output the answer inside <answer></answer>.
Appendix D Visualization of Forecasting Results
To provide a clear comparison among different models, we list supplementary prediction showcases of one representative dataset 8, which are given by the following models: TimeReasoner, PatchTST, DLinear, FEDFormer, iTransformer, AutoFormer, TimeXer, Time-LLM and LLMTime. Among the various models, TimeReasoner predicts the most precise future series variations and exhibits superior performance.
Figure 8: Visualization results of TimeReasoner and other baseline models. It can be seen that TimeReasoner exhibits superior performance.
Appendix E Broader Impacts
This work explores a novel reframing of time series forecasting (TSF) as a conditional reasoning task, leveraging the slow-thinking capabilities of large language models (LLMs). Our findings suggest that LLMs, when guided by structured prompts, can exhibit non-trivial zero-shot forecasting abilities across diverse domains. This paradigm shift has several potential positive and negative societal implications. On the positive side, the TimeReasoner framework may enable more transparent and generalizable forecasting systems, especially in high-stakes domains such as healthcare, finance, and climate modeling. By generating interpretable reasoning traces, the model may serve as a foundation for explainable AI in time-dependent decision-making systems. Furthermore, its ability to function in zero-shot settings could reduce the reliance on task-specific training data, making forecasting more accessible in data-sparse environments or developing regions.
However, the use of general-purpose LLMs in forecasting tasks also raises risks. These models may propagate biases present in the training data, which could lead to unfair or unreliable predictions in critical applications. Additionally, the forecasting performance of LLMs remains sensitive to prompt design and inference-time parameters, which may result in inconsistent outputs or failure under distribution shift. We also acknowledge the substantial computational cost associated with large-scale LLM inference, which may limit its sustainability and accessibility.
From a misuse perspective, the ability to generate plausible but incorrect forecasts could be weaponized in domains such as financial manipulation or disinformation. While our model is intended for research purposes, downstream applications must include appropriate safeguards, such as uncertainty estimation, interpretability tools, and human-in-the-loop oversight.
We encourage future work to examine the ethical implications of deploying LLM-based forecasting systems, develop robust benchmarks for reasoning quality, and explore sustainable approaches to scaling such models. As foundational work in this space, our study is intended to provoke dialogue on responsible AI development and the evolving role of language models in decision-making.

### v2

@@
 AutoFormer MSE 7.4310 13.4821 15.3155 7.0926 0.0012 19662.2860 23335.7346 1659.3374 0.0008 76.7812
-MSE 1.5719 2.2928 2.0689 1.7454 0.0227 47.7715 50.7481 19.3377 0.0226 7.2136
+MAE 1.5719 2.2928 2.0689 1.7454 0.0227 47.7715 50.7481 19.3377 0.0226 7.2136
 FEDFormer MSE 16.1369 13.0165 22.1313 19.1682 0.0013 29623.8156 19774.4204 2168.1601 0.0009 82.3565
-MSE 2.2300 2.3180 2.4822 2.7404 0.0244 55.5783 51.1552 23.9338 0.0215 7.6731
+MAE 2.2300 2.3180 2.4822 2.7404 0.0244 55.5783 51.1552 23.9338 0.0215 7.6731
 PatchTST MSE 9.2430 10.9489 16.3988 5.6375 0.0009 12528.5571 16747.3794 2013.9736 0.0007 50.7402
-MSE 1.6500 1.9930 1.9294 1.3834 0.0198 39.8232 42.0165 20.1837 0.0211 5.6983
+MAE 1.6500 1.9930 1.9294 1.3834 0.0198 39.8232 42.0165 20.1837 0.0211 5.6983
 iTransformer MSE 7.5195 9.9716 12.7432 5.7392 0.0010 13577.9277 17654.6335 1600.0376 0.0008 79.2864
-MSE 1.4983 1.9010 1.6374 1.4422 0.0202 39.8069 41.8849 17.9244 0.0235 7.3322
+MAE 1.4983 1.9010 1.6374 1.4422 0.0202 39.8069 41.8849 17.9244 0.0235 7.3322
 TimeXer MSE 8.4878 11.4123 14.0473 5.5801 0.0009 14509.0913 16505.3908 1668.6824 0.0007 65.8567
-MSE 1.5430 2.0606 1.7534 1.4238 0.0193 41.2693 40.9784 18.2386 0.0230 6.1253
+MAE 1.5430 2.0606 1.7534 1.4238 0.0193 41.2693 40.9784 18.2386 0.0230 6.1253
 DLinear MSE 7.7656 10.3584 14.0352 7.8604 0.0014 21090.0523 21077.8247 1617.3343 0.0007 68.2254
-MSE 1.4981 2.0567 1.7591 1.8981 0.0253 51.1701 50.5644 18.3691 0.0215 6.7148
+MAE 1.4981 2.0567 1.7591 1.8981 0.0253 51.1701 50.5644 18.3691 0.0215 6.7148
 GPT4TS MSE 6.9454 9.7156 15.9028 5.6327 0.0009 13543.6153 16821.9321 1786.3085 0.0010 65.3715
-MSE 1.4127 1.8891 1.9364 1.4743 0.0200 39.7527 41.8708 18.2427 0.0253 6.7207
+MAE 1.4127 1.8891 1.9364 1.4743 0.0200 39.7527 41.8708 18.2427 0.0253 6.7207
 Time-LLM MSE 6.8152 9.9876 15.8078 5.6787 0.0010 13405.9845 16678.0982 1773.9651 0.0011 65.9761
-MSE 1.4115 1.8880 1.9350 1.4733 0.0199 39.7481 41.8598 18.2365 0.0252 6.7151
+MAE 1.4115 1.8880 1.9350 1.4733 0.0199 39.7481 41.8598 18.2365 0.0252 6.7151
 LLMTime MSE 10.6261 11.2011 14.6394 6.9001 0.0026 29141.5143 30266.3043 3979.7996 0.0021 99.5761
-MSE 1.6987 1.5876 1.9543 1.5109 0.0345 65.6789 60.3456 30.1234 0.0234 8.9054
+MAE 1.6987 1.5876 1.9543 1.5109 0.0345 65.6789 60.3456 30.1234 0.0234 8.9054
 TimeReasoner MSE 5.4469 8.6020 14.4327 6.4384 0.0009 11305.6345 12874.9412 1556.5316 0.0008 79.4890
-MSE 1.1984 1.6811 1.7179 1.4209 0.0168 36.1614 37.0193 17.5034 0.0215 6.9735
+MAE 1.1984 1.6811 1.7179 1.4209 0.0168 36.1614 37.0193 17.5034 0.0215 6.9735
 6 Evaluation of TimeReasoner in Different TSF Settings (Q2)
@@
 A.1 Time Series Forecasting
-Time series forecasting has been extensively studied over the past few decades cheng2025comprehensive .
-Early methods, such as ARIMA box1968some ; zhang2003time , employed an auto-regressive approach and forecasted using a moving average framework.
-While theoretical guarantees were established, these traditional methods often require data with ideal properties, which do not always align with real-world scenarios hewamalage2021recurrent .
-With the growing availability of data and advances in computing power, numerous deep learning-based models have emerged, typically following a sequence-to-sequence paradigm salinas2020deepar .
-Recurrent neural networks were among the first to capture temporal dependencies by summarizing past information in time series data wen2017multi .
-However, these architectures suffer from limitations such as a restricted receptive field and error accumulation due to recursive inference lai2018modeling , which undermine forecasting accuracy.
-To improve performance, advanced architectures have been developed to capture long-range dependencies zhou2021informer , including those with self-attention mechanisms and convolutional networks li2019enhancing ; liu2022scinet .
-Recent work has also integrated traditional techniques, like trend-seasonal decomposition and time-frequency transformations zhou2022fedformer ; liu2024generative , into neural networks to better capture time series data characteristics wen2020fast ; zhou2022film ; cheng2025convtimenet .
-Additionally, a recent study showed that even a simple linear network enhanced with decomposition can achieve competitive performance zeng2023transformers .
-Furthermore, slice-based methods have demonstrated superior accuracy in long-term forecasting tasks nie2022time ; zhang2023crossformer ; cheng2025instructime .
+Time series forecasting has been extensively studied over the past few decades [11].
+Early methods, such as ARIMA [3, 76], employed an auto-regressive approach and forecasted using a moving average framework.
+While theoretical guarantees were established, these traditional methods often require data with ideal properties, which do not always align with real-world scenarios [22].
+With the growing availability of data and advances in computing power, numerous deep learning-based models have emerged, typically following a sequence-to-sequence paradigm [55].
+Recurrent neural networks were among the first to capture temporal dependencies by summarizing past information in time series data [67].
+However, these architectures suffer from limitations such as a restricted receptive field and error accumulation due to recursive inference [33], which undermine forecasting accuracy.
+To improve performance, advanced architectures have been developed to capture long-range dependencies [82], including those with self-attention mechanisms and convolutional networks [37, 43].
+Recent work has also integrated traditional techniques, like trend-seasonal decomposition and time-frequency transformations [84, 47], into neural networks to better capture time series data characteristics [66, 83, 12].
+Additionally, a recent study showed that even a simple linear network enhanced with decomposition can achieve competitive performance [73].
+Furthermore, slice-based methods have demonstrated superior accuracy in long-term forecasting tasks [50, 81, 10].
 A.2 Large Language Models for Time Series
-In recent years, large language models (LLMs) have garnered significant attention for their remarkable ability to understand and generate human-like text kumar2024large ; yao2024survey .
-While traditionally applied to natural language processing tasks zhang2024large , these models have also shown great potential in time series analysis  jin2024position .
-The application of LLMs to time series can be broadly categorized into two primary paradigms: fine-tuning and prompt-based zero-shot learning jiang2024empowering .
-In fine-tuning, a pre-trained LLM is further trained on time series data specific to a given task chang2023llm4ts ; chang2024align , enabling it to learn task-specific patterns and improve performance  liu2025calf ; jin2023time .
-However, this approach often requires substantial labeled data and computational resources cheng2025comprehensive .
-On the other hand, prompt-based zero-shot methods leverage the model’s pre-existing knowledge gruver2023large , guiding its predictions through well-crafted prompts without the need for task-specific training liu2024lstprompt .
-Although zero-shot methods are more flexible and resource-efficient tang2025time , they may not always achieve the same level of performance as fine-tuned models, especially on specialized time series tasks wang2024tabletime .
+In recent years, large language models (LLMs) have garnered significant attention for their remarkable ability to understand and generate human-like text [32, 71].
+While traditionally applied to natural language processing tasks [80], these models have also shown great potential in time series analysis  [28].
+The application of LLMs to time series can be broadly categorized into two primary paradigms: fine-tuning and prompt-based zero-shot learning [26].
+In fine-tuning, a pre-trained LLM is further trained on time series data specific to a given task [4, 5], enabling it to learn task-specific patterns and improve performance  [44, 27].
+However, this approach often requires substantial labeled data and computational resources [11].
+On the other hand, prompt-based zero-shot methods leverage the model’s pre-existing knowledge [19], guiding its predictions through well-crafted prompts without the need for task-specific training [41].
+Although zero-shot methods are more flexible and resource-efficient [59], they may not always achieve the same level of performance as fine-tuned models, especially on specialized time series tasks [63].
 Both paradigms highlight the growing interest in applying LLMs to time series, although challenges remain in optimizing effectiveness for these tasks.
 A.3 Inference-Time Reasoning
-Inference-Time Reasoning zeng2025revisiting refers to the strategy of enhancing a pre-trained model’s performance on complex reasoning tasks by allocating additional computational resources during the inference phase singhi2025solve ; yang2025towards .
-These resources may include extended inference time, more computational steps, or advanced reasoning strategies such as multi-path sampling and iterative refinement leehanchung .
-Unlike traditional training-time scaling, which focuses on improving a model’s capabilities during training through increased parameters, data, or computation parashar2025inference , inference-time reasoning utilizes a larger computational budget at inference to unlock deeper reasoning potential within the model—potential zhang2025and that might otherwise remain dormant during rapid inference.
-This approach improves performance without altering the model’s weights chen2024simple .
-Several cutting-edge large language models demonstrate the effectiveness of inference time reasoning sui2025meta .
-For instance, DeepSeek-R1 guo2025deepseek can benefit from techniques such as iterative refinement or self-correction, utilizing additional computational resources during inference to enhance performance on complex tasks, leveraging its strengths in coding and reasoning mercer2025brief .
-Similarly, models such as Gemini-2.5-pro team2023gemini are expected to excel in sophisticated reasoning tasks when granted extended processing time or multi-step inference protocols enabled by inference-time reasoning.
-OpenAI’s models, including versions such as GPT-o1 hasei2025comparative , show significant improvements on challenging reasoning problems when employing techniques such as chain-of-thought prompting or step-by-step decomposition, both of which increase the computational budget during inference wei2022chain .
-These examples highlight the growing importance of inference-time reasoning in pushing the boundaries of AI reasoning by trading inference-time computation for higher accuracy muennighoff2025s1 .
+Inference-Time Reasoning [75] refers to the strategy of enhancing a pre-trained model’s performance on complex reasoning tasks by allocating additional computational resources during the inference phase [56, 70].
+These resources may include extended inference time, more computational steps, or advanced reasoning strategies such as multi-path sampling and iterative refinement [35].
+Unlike traditional training-time scaling, which focuses on improving a model’s capabilities during training through increased parameters, data, or computation [51], inference-time reasoning utilizes a larger computational budget at inference to unlock deeper reasoning potential within the model—potential [77] that might otherwise remain dormant during rapid inference.
+This approach improves performance without altering the model’s weights [9].
+Several cutting-edge large language models demonstrate the effectiveness of inference time reasoning [57].
+For instance, DeepSeek-R1 [20] can benefit from techniques such as iterative refinement or self-correction, utilizing additional computational resources during inference to enhance performance on complex tasks, leveraging its strengths in coding and reasoning [48].
+Similarly, models such as Gemini-2.5-pro [61] are expected to excel in sophisticated reasoning tasks when granted extended processing time or multi-step inference protocols enabled by inference-time reasoning.
+OpenAI’s models, including versions such as GPT-o1 [21], show significant improvements on challenging reasoning problems when employing techniques such as chain-of-thought prompting or step-by-step decomposition, both of which increase the computational budget during inference [65].
+These examples highlight the growing importance of inference-time reasoning in pushing the boundaries of AI reasoning by trading inference-time computation for higher accuracy [49].
 Appendix B Experimental Details
@@
 B.2 Baselines
-  • Autoformer wu2021autoformer : Introduces a decomposition-based Transformer architecture with an auto-correlation mechanism to capture long-term dependencies in time series forecasting.
-  • FEDformer zhou2022fedformer : Combines seasonal-trend decomposition with frequency domain analysis using Fourier transforms to enhance long-term time series forecasting.
-  • PatchTST nie2022time : Segments time series into patches and applies a channel-independent Transformer to improve long-term forecasting accuracy.
-  • iTransformer liu2023itransformer : Applies attention mechanisms on inverted dimensions, embedding time points as variate tokens to capture multivariate correlations in time series data.
-  • TimeXer wang2024timexer : Integrates exogenous variables into Transformer-based time series forecasting, enhancing predictions by incorporating external information.
-  • DLinear zeng2023transformers : A simple and efficient linear model that decomposes time series into trend and seasonal components, achieving competitive performance with reduced complexity.
-  • GPT4TS zhou2023one : Utilizes pre-trained GPT-2 models for time series forecasting by feeding time series patches in a channel-independent manner.
-  • Time-LLM jin2023time : Reprograms large language models for time series forecasting by converting time series data into textual prompts, enabling few-shot and zero-shot learning capabilities.
-  • LLMTime gruver2023large : Employs large language models for zero-shot time series forecasting by encoding numerical data as text and generating predictions through text completions.
+  • Autoformer [68]: Introduces a decomposition-based Transformer architecture with an auto-correlation mechanism to capture long-term dependencies in time series forecasting.
+  • FEDformer [84]: Combines seasonal-trend decomposition with frequency domain analysis using Fourier transforms to enhance long-term time series forecasting.
+  • PatchTST [50]: Segments time series into patches and applies a channel-independent Transformer to improve long-term forecasting accuracy.
+  • iTransformer [45]: Applies attention mechanisms on inverted dimensions, embedding time points as variate tokens to capture multivariate correlations in time series data.
+  • TimeXer [64]: Integrates exogenous variables into Transformer-based time series forecasting, enhancing predictions by incorporating external information.
+  • DLinear [73]: A simple and efficient linear model that decomposes time series into trend and seasonal components, achieving competitive performance with reduced complexity.
+  • GPT4TS [85]: Utilizes pre-trained GPT-2 models for time series forecasting by feeding time series patches in a channel-independent manner.
+  • Time-LLM [27]: Reprograms large language models for time series forecasting by converting time series data into textual prompts, enabling few-shot and zero-shot learning capabilities.
+  • LLMTime [19]: Employs large language models for zero-shot time series forecasting by encoding numerical data as text and generating predictions through text completions.
 B.3 Robustness Evaluation
@@
 Normalization is a key preprocessing step in time series forecasting.
-We compare two strategies: reversible instance normalization (RevIN) kim2021reversible and Z-score normalization, which differ in how they compute and apply normalization statistics.
+We compare two strategies: reversible instance normalization (RevIN) [30] and Z-score normalization, which differ in how they compute and apply normalization statistics.
 Reversible Instance Normalization.